# Coupled Axial and Transverse Currents Method for Finite Element Modelling of Periodic Superconductors

### v2

@@ 3.1 Reference model
We consider the same 2D geometry and mesh of the cross-section for the CATI and helicoidal methods. The chosen mesh leads to a number of 63k DOFs with the CATI method and 113k DOFs with the helicoidal method.
Note that the two models do not define exactly the same physical problem. Indeed, the helicoidal model describes tilted filaments rigorously, with three-dimensional current flow, whereas this is neglected with the CATI method, as the filaments are modelled as straight conductors. As a result, the filaments in the helicoidal method have a reduced effective cross-section compared to the circular section in the $x$-$y$-plane. For real strands however the importance of this effect is limited. The effective cross-section reduction of a filament is of the order of $1-\cos\theta$ with $\theta$ the tilt angle of that filament with respect to the $\hat{\boldsymbol{e}}_{z}$ direction. For the typical value $p/d=20$, $\theta<9\degree$ and a maximum difference of the order of $1.2$% is therefore expected locally on the geometry.
@@ 3.2 Results and error map
Figure 6: Total AC loss per cycle and separate contributions versus the frequency $f$ of an applied transverse magnetic field (of amplitude $1$ T), for $p/d=20$ and the 54-filament strand with linear material properties. Solid curves are for the CATI method. Dashed curves are for the reference helicoidal method.
@@
Figure 7: Energy loss density per cycle $q_{\text{eddy}}+q_{\text{coupling}}$ in the matrix for $p/d=20$ with linear material properties, at four different frequencies. Each case is represented on one quarter of the cross-section, dash-dotted lines represent symmetry lines. The colormap is in a log-scale.
@@
Figure 8: Total AC loss per cycle and per unit length as a function of the frequency $f$ of an applied transverse magnetic field, for different twist pitch lengths to diameter ratios $p/d$ for the 54-filament strand with linear material properties. Solid curves are for the CATI method. Dashed curves are for the reference helicoidal method.
@@ 3.3 Discussion of the length correction factor
Figure 10: Effect of the length correction factor. Solid curves are results of the CATI method without the length correction, i.e., using $\ell$ instead of $\ell^{\text{c}}$. Dashed curves are reference results of the helicoidal method.
@@ 3.4 Correction for dynamic effects in the matrix
Figure 12: Power loss density $p_{\text{coupling}}$ due to transverse coupling currents in harmonic regime for $p/d=20$ in four different cases with linear material properties. Each case is represented on one quarter of the cross-section, dash-dotted lines represent symmetry lines. The colormap is in a log-scale. (a-b) Reference helicoidal solutions on the plane $z=0$, at $f=100$ Hz and $f=1$ kHz, respectively. (c-d) CATI method solutions at $f=1$ kHz, without and with the dynamic correction, respectively.
@@ 4 Verification with Nonlinear Materials
$\rho_{\text{SC}}(\boldsymbol{j},\boldsymbol{b})=\frac{e_{\text{c}}}{j_{\text{c}}(\boldsymbol{b})}\left(\frac{\|\boldsymbol{j}\|}{j_{\text{c}}(\|\boldsymbol{b}\|)}\right)^{n-1},$ (25)
@@ 5.1 Loss map and magnetization curves
As a first application of the CATI method, we model the 54-filament strand (see Table 1) response to an external transverse magnetic field, varying sinusoidally with frequencies $f$ ranging from $0.01$ Hz to $10$ kHz, sampled with $25$ values, and amplitudes $b$ from $0.001$ T to $5.6$ T, sampled with $16$ values. We fix the twist pitch length to $p=19$ mm. A loss map, inspired by the work of A. M. Campbell in [12], is plotted in Fig. 15 and gives the total AC loss per cycle and per unit length of the strand, interpolated from the results ($400$ simulations) in this parameter space. The AC loss per cycle is computed as described in Section 4. The different loss contribution are also illustrated as a function of frequency in Fig. 16 for a field amplitude of $0.2$ T.
Figure 15: Total loss per cycle and per unit length as a function of the frequency and amplitude of an applied transverse field, and no transport current. Dashed curves are contour lines of equal loss. Solid curves delimit areas of distinct dominant loss contributions (filament, coupling, or eddy losses). The four points are at positions where the solution is represented in Fig. 17(a-d). This loss map is inspired from [12].
Figure 16: Total loss per cycle and its loss contributions as a function of the frequency $f$ of an applied transverse magnetic field of amplitude $b=0.2$ T.
@@
The frequency dependence of the coupling and eddy loss contributions in Fig. 16 is comparable to that obtained with linear materials in Section 3. The coupling loss follows a typical bell curve, approached by analytical models by Eq. (20[12], and the shape of the eddy loss curve is associated with skin effect in the matrix.
By contrast, the filament loss curve of superconducting (nonlinear) filaments is qualitatively different from that of resistive (linear) filaments. In the nonlinear case, the evolution of the filament loss with frequency can be decomposed in four regimes, see Fig. 17(a-d). First, at the lowest frequencies, filaments are fully uncoupled and exhibit hysteresis losses that are, per cycle, almost frequency-independent. As the frequency further increases, coupling currents start to progressively couple the filaments, which leads to a smooth transition of the filament loss per cycle towards a new plateau, during which the coupling loss dominates the total loss. When filaments are fully coupled, the filament loss per cycle is again almost frequency-independent. Whether the filament loss for coupled filaments is larger or smaller than the filament loss for uncoupled filaments depends on the external field amplitude. Finally, at the highest frequencies, eddy currents in the matrix start to shield the inner part of the strand, leading to a sharp decrease of filament losses, because the field seen by the superconducting filaments is decreasing. In this regime, eddy current loss dominates the total loss.
@@
The ability of the CATI method to model many different configurations in a reduced amount of time is useful in the context of homogenization methods. The idea for such methods is to replace detailed geometries, e.g., a cable made of many multifilamentary strands, or a magnet made up of many such cables, by simplified geometries with appropriate homogenized material parameters reproducing the macroscopic response of the detailed models without describing the details of them explicitly [47, 48]. The identification of these material parameters requires the knowledge of the detailed models response under a range of different excitations. The CATI method provides an efficient way of computing this accurately.
@@ 6 Conclusion
In this contribution, we propose the Coupled Axial and Transverse currents (I) (CATI) method. It is based on a pair of coupled 2D models for modelling transient electric and magnetic effects in composite superconducting conductors with a periodic geometry. We applied the CATI method to twisted superconducting strands and verified its accuracy by comparison against reference methods. We demonstrated that the CATI method reproduces the coupling current density and all types of losses with a very good accuracy in a wide range of frequencies, in the case where the twist pitch length is large enough compared to the strand diameter ($p/d\gtrsim 10$).
As the CATI method is based on 2D models, it leads to a massive reduction of the computational cost compared to conventional 3D models. In particular, this allows to perform detailed parameter sweep studies and in-depth analyses of the strand response in many different configurations in a reasonable amount of time. As an application, such parameter studies could be exploited in the context of homogenization methods, if they are used as detailed mesoscale models for identifying equivalent material parameters of homogenized macroscale models [47].
We presented the practical implementation of the CATI method, and highlighted the ease of it, as any FE software allowing for field-circuit coupling is in principle ready for its implementation. The use of circuit equations for coupling the AI and TI models also allows for a direct extension of the method for the treatment of contact resistance between the filament and the matrix, as well as that of diffusion barriers in Nb3Sn wires [49], that may be modelled as thin-shells [50].
Compared to the helicoidal transformation technique, in addition of handling much more easily nonlinear materials, the CATI method also applies to non-helicoidally symmetric cross-sections, such as deformed geometries [49]. Furthermore, an extension of the approach to other types of conductors with periodic geometry, e.g., cables with transposed conductors such as Roebel cables [33], or Rutherford cables [32], should be straight-forward, and will be subject of further research.

### v3

@@ 3.1 Reference model
We consider the same 2D geometry and mesh of the cross section for the CATI and helicoidal methods. The chosen mesh leads to a number of 63k DOFs with the CATI method and 113k DOFs with the helicoidal method.
Note that the two models do not define exactly the same physical problem. Indeed, the helicoidal model describes tilted filaments rigorously, with three-dimensional current flow, whereas this is neglected with the CATI method, as the filaments are modelled as straight conductors. As a result, the filaments in the helicoidal method have a reduced effective cross section compared to the circular section in the $x$-$y$-plane. For real strands however the importance of this effect is limited. The effective cross section reduction of a filament is of the order of $1-\cos\theta$ with $\theta$ the tilt angle of that filament with respect to the $\hat{\boldsymbol{e}}_{z}$ direction. For the typical value $p/d=20$, $\theta<9\degree$ and a maximum difference of the order of $1.2$% is therefore expected locally on the geometry.
@@ 3.2 Results and error map
Figure 6: Total AC loss per cycle and separate contributions versus the frequency $f$ of an applied transverse magnetic field (of amplitude $1$ T), for $p/d=20$ and the 54-filament strand with linear material properties. Solid curves are results of the CATI method. Dashed curves are results of the reference helicoidal method.
@@
Figure 7: Energy loss density per cycle $q_{\text{eddy}}+q_{\text{coupling}}$ in the matrix for $p/d=20$ with linear material properties, at four different frequencies. Each case is represented on one quarter of the cross section, dash-dotted lines represent symmetry lines. The colormap is in a log-scale.
@@
Figure 8: Total AC loss per cycle and per unit length as a function of the frequency $f$ of an applied transverse magnetic field, for different twist pitch lengths to diameter ratios $p/d$ for the 54-filament strand with linear material properties. Solid curves are results of the CATI method. Dashed curves are results of the reference helicoidal method.
@@ 3.3 Discussion of the length correction factor
Figure 10: Effect of the length correction factor. Solid curves are results of the CATI method without the length correction, i.e., using $\ell$ and not $\ell^{\text{c}}$ in Eqn. (2.1.1) and (9). Dashed curves are reference results of the helicoidal method.
@@ 3.4 Correction for dynamic effects in the matrix
Figure 12: Power loss density $p_{\text{coupling}}$ due to transverse coupling currents in harmonic regime for $p/d=20$ in four different cases with linear material properties. Each case is represented on one quarter of the cross section, dash-dotted lines represent symmetry lines. The colormap is in a log-scale. (a-b) Reference helicoidal solutions on the plane $z=0$, at $f=100$ Hz and $f=1$ kHz, respectively. (c-d) CATI method solutions at $f=1$ kHz, without and with the dynamic correction, respectively.
@@ 4 Verification with Nonlinear Materials
$\rho_{\text{SC}}(\boldsymbol{j},\boldsymbol{b})=\frac{e_{\text{c}}}{j_{\text{c}}(\|\boldsymbol{b}\|)}\left(\frac{\|\boldsymbol{j}\|}{j_{\text{c}}(\|\boldsymbol{b}\|)}\right)^{n-1},$ (25)
@@ 5.1 Loss map and magnetization curves
As a first application of the CATI method, we consider the 54-filament strand (see Table 1) with the same material properties as in Section 4. We model its response to an external transverse magnetic field, varying sinusoidally with frequencies $f$ ranging from $0.01$ Hz to $10$ kHz, sampled with $25$ values, and amplitudes $b$ from $0.001$ T to $5.6$ T, sampled with $16$ values. We fix the twist pitch length to $p=19$ mm. A loss map, inspired by the work of A. M. Campbell in [12], is plotted in Fig. 15 and gives the total AC loss per cycle and per unit length of the strand, interpolated from the results ($400$ simulations) in this parameter space. The AC loss per cycle is computed as described in Section 4. The different loss contributions are also illustrated as a function of frequency in Fig. 16 for field amplitudes of $0.02$ T, $0.2$ T, and $2$ T.
Figure 15: Total loss per cycle and per unit length as a function of the frequency and amplitude of an applied transverse field, and no transport current. Dashed curves are contour lines of equal loss. Solid curves delimit areas of distinct dominant loss contributions (filament, coupling, or eddy losses). The four points are at positions where the solution is represented in Fig. 17(a-d). This loss map is inspired from the work of A. M. Campbell in [12].
$10^{-2}$$10^{-1}$$10^{0}$$b=2$ TAC pcy(J/m)
$10^{-2}$$10^{-1}$$10^{0}$$10^{-4}$$10^{-3}$$10^{-2}$$b=0.2$ TAC Loss per cycle (J/m)TotalFilamentCouplingEddy
$10^{-2}$$10^{-1}$$10^{0}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{-6}$$10^{-5}$$10^{-4}$$b=0.02$ TFrequency $f$ (Hz)AC pcy(J/m)TotalFilamentCouplingEddy
Figure 16: Total loss per cycle and separated loss contributions as a function of the frequency $f$, for three different field amplitudes. The legend is the same for the three subfigures.
@@
The frequency dependence of the coupling and eddy loss contributions in Fig. 16 is comparable to that obtained with linear materials in Section 3. The coupling loss follows a typical bell curve [12], and the shape of the eddy loss curve is associated with skin effect in the matrix. A good match with the analytical model is obtained at low field amplitudes ($b\lesssim 0.1$ T). At higher field amplitudes, saturation effects in the filaments affect the dynamics, which is no longer well described by the analytical model.
The filament loss curve of superconducting (nonlinear) filaments is qualitatively different from that of resistive (linear) filaments. In the nonlinear case, the evolution of the filament loss with frequency can be decomposed in four regimes, see Fig. 17(a-d). First, at the lowest frequencies, filaments are fully uncoupled and exhibit hysteresis losses that are, per cycle, almost frequency-independent. As the frequency further increases, coupling currents start to progressively couple the filaments, which leads to a smooth transition of the filament loss per cycle towards a new plateau, during which the coupling loss dominates the total loss. When filaments are fully coupled, the filament loss per cycle is again almost frequency-independent. Whether the filament loss for coupled filaments is larger or smaller than the filament loss for uncoupled filaments depends on the external field amplitude. Finally, at the highest frequencies, eddy currents in the matrix start to shield the inner part of the strand, leading to a sharp decrease of filament losses, because the field seen by the superconducting filaments is decreasing. In this regime, eddy current loss dominates the total loss.
@@
The ability of the CATI method to model many different configurations in a reduced amount of time is useful in the context of homogenization methods. The idea for such methods is to replace detailed geometries, e.g., a cable made of many multifilamentary strands, or a magnet made of many such cables, by simplified geometries with appropriate homogenized material parameters reproducing the macroscopic response of the detailed models without describing the details of them explicitly [48, 49]. The identification of these material parameters requires the knowledge of the detailed models response under a range of different excitations. The CATI method provides an efficient way of computing this accurately.
@@ 6 Conclusion
In this contribution, we proposed the Coupled Axial and Transverse currents (I) (CATI) method. It is based on a pair of coupled 2D models for modelling transient electric and magnetic effects in composite superconducting conductors with a periodic geometry. We applied the CATI method to twisted superconducting strands and verified its accuracy by comparison against reference methods. We demonstrated that the CATI method reproduces the coupling current density and all types of losses with a very good accuracy in a wide range of frequencies, in the case where the twist pitch length is large enough compared to the strand diameter ($p/d\gtrsim 10$).
As the CATI method is based on 2D models, it leads to a massive reduction of the computational cost compared to conventional 3D models, from more than a week (168h) to less than an hour for the tested cases. In particular, this allows to perform detailed parameter sweep studies and in-depth analyses of the strand response in many different configurations in a reasonable amount of time, and with a reasonable amount of computational resources. As an application, such parameter studies could be exploited in the context of homogenization methods, if they are used as detailed mesoscale models for identifying equivalent material parameters of homogenized macroscale models [48].
We presented the practical implementation of the CATI method, and highlighted the ease of it. Any FE software allowing for field-circuit coupling is in principle ready for its implementation. The use of circuit equations for coupling the AI and TI models also allows for a direct extension of the method for the treatment of contact resistance between the filament and the matrix, as well as that of diffusion barriers in Nb3Sn wires [50], that may be modelled as thin-shells [51].
Compared to the helicoidal transformation technique, in addition of handling much more easily nonlinear materials, the CATI method also applies to non-helicoidally symmetric cross sections, such as deformed geometries [50]. Furthermore, an extension of the approach to other types of conductors with periodic geometry, e.g., cables with transposed conductors such as Roebel cables [34], or Rutherford cables [33], will be subject of further research.